OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. A(4, 1), B(5, 2)

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

skope-intersept (1-2)/(4-5)=1 y=x-3

OpenStudy (alexandervonhumboldt2):

x-y=3

OpenStudy (alexandervonhumboldt2):

standart x-y=3

OpenStudy (anonymous):

thnx